OpenStudy (anonymous):

6-3(2x-4)=2(1+x) solve for X

OpenStudy (anonymous):

6-3(2x-4)=2(1+x) 6-6x+12=2+2x 18-6x=2+2x 16-6x=2x 16=8x x=2